# Taylor-Accelerated Neural Network Interpolation Operators on Irregular Grids with Higher Order Approximation

### v1

@@ Abstract
In this paper, a new class of Taylor-accelerated neural network interpolation operators is introduced on quasi-uniform irregular grids. These operators improve existing neural network interpolation operators by incorporating Taylor polynomials at the sampling nodes, thereby exploiting higher smoothness of the target function. The proposed operators are shown to be well defined, uniformly bounded, and to satisfy an exact interpolation property at the grid points. In addition, polynomial reproduction up to a prescribed degree is established. Direct approximation estimates are derived in terms of higher-order moduli of smoothness, yielding enhanced convergence rates for sufficiently smooth functions. Numerical experiments are presented to support the theoretical analysis and to demonstrate the significant accuracy improvement achieved through the Taylor-accelerated construction. In particular, higher-order convergence on irregular grids is obtained, and the proposed approach outperforms existing neural network interpolation operators on irregular partitions, including Lagrange-based schemes.
@@ 1 Introduction and Preliminaries
The systematic study of NNOs was initiated by Cardaliaguet and Euvrard [4], who introduced bell-shaped and squashing-type neural network operators and established early connections between neural network structures and polynomial-type approximation processes. Subsequently, Anastassiou and his collaborators [1, 2, 3] developed a broad approximation theory for sigmoidal-based neural network operators, including multivariate extensions and convergence rates in various normed settings. Afterward, several authors investigated NNOs using different activation functions and normed spaces for the functions under approximation [8, 11, 5, 6, 17].
In recent years, increasing attention has been devoted to interpolation-type neural network operators, which aim to reconstruct functions directly from sampled data. One of the first constructive interpolation operators within the NNOs framework was proposed by Costarelli [9], and later extended to multivariate settings in [10]. Further developments include interpolation operators activated by smooth ramp functions, providing quantitative rates in terms of moduli of continuity and smoothness (see, e.g., [16, 19, 14, 15]).
On the other hand, existing neural network interpolation operators often achieve only low-order approximation, since their error is typically governed by the first modulus of continuity even when the target function possesses additional smoothness. This motivates the problem of designing neural network interpolation operator capable of exploiting higher regularity and attaining improved convergence orders.
@@
Inspired by higher-order sampling methods and local polynomial approximation techniques, the present work introduces a new class of Taylor-accelerated neural network interpolation operators on irregular partitions. The main idea is to enrich sampling-based neural network interpolation operators by replacing pointwise function values with local Taylor polynomials of degree $r$, thereby incorporating derivative information into the approximation mechanism. As a result, the proposed operators preserve exact interpolation at the grid nodes while achieving approximation order $r+1$ for sufficiently smooth functions, with error estimates expressed in terms of higher-order moduli of smoothness.
Now, we collect some definitions and tools that will be used throughout the paper. In particular, we recall the functional setting, the notion of irregular partitions, the class of compactly supported sigmoidal activation, and higher-order moduli of smoothness.
@@
Since the Taylor-accelerated neural network operators introduced in this work employ local Taylor polynomials of degree $r$, we shall mainly assume $g\in C^{r}[c,d]$. Moreover, in the derivation of quantitative error estimates we will require the additional smoothness condition $g\in C^{r+1}[c,d]$.
Irregular Partitions. Let $c=z_{0}<z_{1}<\cdots<z_{n}=d$ be a partition of the interval $[c,d]$. In contrast to uniform grids, the nodes $\{z_{k}\}_{k=0}^{n}$ are allowed to be distributed in a non-uniform manner.
@@
    Indeed, since the partition $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{k+1}-z_{k}\geq d_{1}$ for all $k$, it follows that
@@ 2 Taylor-Accelerated Neural Network Interpolation Operator on Irregular Grids
In this section, we introduce a Taylor-type neural network interpolation operator constructed on irregular partitions. The main idea is to enhance a neural network interpolation operator by incorporating higher-order Taylor information at the grid nodes, while preserving the exact nodal interpolation property.
@@
Let $g\in C^{r}[c,d]$. The Taylor-accelerated neural network interpolation operator is defined by
@@
where $\sigma$ is the compactly supported activation function introduced in Section 1 and $d_{1}$ denotes the minimum mesh size of the irregular partition.
@@ 2.1 Approximation Estimate
Since $\deg(p)\leq r$, we have $p^{(r+1)}\equiv 0$. Therefore, the Taylor polynomial of degree $r$ of $p$ about any node $z_{k}$ coincides with $p$ itself. Hence, for all $y\in[c,d]$ and all $k=0,1,\dots,n$,
@@
Next, if $k\neq j$, then since the partition $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{i+1}-z_{i}\geq d_{1}$ for all $i$, we have
@@
denotes the mesh norm of the irregular partition. In particular, if $d_{2}\to 0$, then $\widetilde{T}_{n,r}(g)\to g$ uniformly on $[c,d]$.
@@
denote the mesh norm of the irregular partition. Then the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$ satisfies
@@
Assume that the sequence of irregular partitions satisfies the refinement condition: there exists a constant $C>0$, independent of $n$, such that
@@
Such a condition is fulfilled, for instance, by any quasi-uniform family of partitions of $[c,d]$.
@@
Although the operator introduced in [7] is defined for all continuous functions $g\in C[c,d]$, its higher-order accuracy is obtained through a Lagrange-type reconstruction involving $(r+1)$ neighboring sampling nodes. On irregular partitions, this may lead to additional numerical complexity and reduced stability due to the behaviour of Lagrange basis polynomials.
@@ 3.1 Activation Function and Grid Construction
All experiments were implemented in MATLAB. The sampling nodes were chosen as nested quasi-uniform irregular partitions of $[0,1]$, generated by applying small random perturbations to a uniform grid while preserving the mesh regularity condition $\frac{d_{2}}{d_{1}}<2.$ This construction ensures that the theoretical assumptions are satisfied and allows us to study convergence as the number of nodes $n$ increases.
@@ 3.2 Error Convergence and Comparison
where $T^{\mathrm{Lag}}_{n,r}$ is the Lagrange-based neural network operator on irregular grids introduced by Costarelli et al.[7].
@@
To highlight the improvement with respect to the Lagrange-based construction, Figures 2-4 compare the Taylor and Lagrange operators for each fixed order $r=1,2,3$. In all cases, the Taylor-accelerated operator provides smaller errors, demonstrating the effectiveness of the local Taylor enrichment.
Figure 2: Error comparison between Taylor and Lagrange neural network operators for $r=1$.
Figure 3: Error comparison between Taylor and Lagrange neural network operators for $r=2$.
Figure 4: Error comparison between Taylor and Lagrange neural network operators for $r=3$.
@@ 4 Conclusion
We proved that the operators are well defined, uniformly bounded, and preserve exact interpolation at the grid nodes. Moreover, they reproduce polynomials up to the prescribed order and satisfy direct approximation estimates in terms of higher-order moduli of smoothness, leading to convergence rates of order $r+1$ for sufficiently smooth functions. In particular, under asymptotically quasi-uniform partitions, the error decays as
@@
Numerical experiments confirmed the theoretical findings and showed that the Taylor-accelerated operator provides improved accuracy compared not only with classical first-order neural network interpolation operators, but also with the recently proposed Lagrange-based neural network interpolation operator on irregular grids. This highlights the advantage of incorporating local Taylor information in achieving higher-order convergence on non-uniform sampling partitions.
Future research directions include multivariate extensions, adaptive irregular meshes, and the development of stochastic or fuzzy variants of the proposed Taylor-accelerated framework.

### v2

@@ Abstract
In this paper, a new class of Taylor-accelerated neural network interpolation operators is introduced on quasi-uniform irregular grids. These operators improve existing neural network interpolation operators by incorporating Taylor polynomials at the sampling nodes, thereby exploiting higher smoothness of the target function. The proposed operators are shown to be well defined, uniformly bounded, and to satisfy an exact interpolation property at the grid points. In addition, polynomial reproduction up to a prescribed degree is established. Jackson-type approximation estimates are derived in terms of higher-order moduli of smoothness, yielding enhanced convergence rates for sufficiently smooth functions. Numerical experiments are presented to support the theoretical analysis and to demonstrate the significant accuracy improvement achieved through the Taylor-accelerated construction. In particular, higher-order convergence on irregular grids is obtained, and the proposed approach outperforms existing neural network interpolation operators on irregular grids, including Lagrange-based schemes.
@@ 1 Introduction and Preliminaries
The systematic study of NNOs was initiated by Cardaliaguet and Euvrard [4], who introduced bell-shaped and squashing-type NNOs and established early connections between neural network structures and polynomial-type approximation processes. Subsequently, Anastassiou and his collaborators [1, 2, 3] developed a broad approximation theory for sigmoidal-based neural network operators, including multivariate extensions and convergence rates in various normed settings. Afterward, several authors investigated NNOs using different activation functions and normed spaces for the functions under approximation [8, 11, 5, 6, 18].
On the other hand, existing NNOs often achieve only low-order approximation, since their error is typically governed by the first modulus of continuity even when the target function possesses additional smoothness. This motivates the problem of designing neural network operator capable of exploiting higher regularity and attaining improved convergence orders. To address this limitation, Kadak[14] recently proposed Hermite-type NNOs incorporating derivative information of the approximated function on uniform grids.
In recent years, increasing attention has been devoted to interpolation-type NNOs, which aim to reconstruct functions directly from sampled data. One of the first constructive interpolation operators within the NNOs framework was proposed by Costarelli [9], and later extended to multivariate settings in [10]. Further developments include interpolation operators activated by smooth ramp functions, providing quantitative rates in terms of moduli of continuity and smoothness (see, e.g., [17, 20, 15, 16]).
@@
Inspired by higher-order sampling methods and local polynomial approximation techniques, the present work introduces a new class of Taylor-accelerated neural network interpolation operators on irregular grids. The main idea is to enrich sampling-based neural network interpolation operators by replacing pointwise function values with local Taylor polynomials of degree $r$, thereby incorporating derivative information into the approximation mechanism. As a result, the proposed operators preserve exact interpolation at the grid nodes while achieving approximation order $r+1$ for sufficiently smooth functions, with error estimates expressed in terms of higher-order moduli of smoothness.
Now, we collect some definitions and tools that will be used throughout the paper. In particular, we recall the functional setting, the notion of irregular grids, the class of compactly supported sigmoidal activation, and higher-order moduli of smoothness.
@@
Since the Taylor-accelerated neural network interpolation operators introduced in this work employ local Taylor polynomials of degree $r$, we shall mainly assume $g\in C^{r}[c,d]$. Moreover, in the derivation of quantitative error estimates we will require the additional smoothness condition $g\in C^{r+1}[c,d]$.
Irregular grids. Let $c=z_{0}<z_{1}<\cdots<z_{n}=d$ be a partition of the interval $[c,d]$. In contrast to uniform grids, the nodes $\{z_{k}\}_{k=0}^{n}$ are allowed to be distributed in a non-uniform manner.
@@
    Indeed, since the grids $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{k+1}-z_{k}\geq d_{1}$ for all $k$, it follows that
@@ 2 Taylor-Accelerated Neural Network Interpolation Operator on Irregular Grids
In this section, we introduce a Taylor-type neural network interpolation operator constructed on irregular grids. The main idea is to enhance a neural network interpolation operator by incorporating higher-order Taylor information at the grid nodes, while preserving the exact nodal interpolation property.
@@
Let $g\in C^{r}[c,d]$. Then Taylor-accelerated neural network interpolation operator is defined by
@@
where $\sigma$ is the compactly supported activation function introduced in Section 1 and $d_{1}$ denotes the minimum mesh size of the irregular grids.
@@ 2.1 Approximation Estimate
Since $\deg(p)\leq r$, we have $p^{(r+1)}\equiv 0$. Therefore, the Taylor polynomial of degree $r$ of $p$ about any node $z_{k}$ coincides with $p$ itself, i.e., for all $y\in[c,d]$ and all $k=0,1,\dots,n$,
@@
Next, if $k\neq j$, then since the grids $\{z_{k}\}_{k=0}^{n}$ is strictly increasing and satisfies $z_{i+1}-z_{i}\geq d_{1}$ for all $i$, we have
@@
denotes the mesh norm of the irregular grids. In particular, if $d_{2}\to 0$, then $\widetilde{T}_{n,r}(g)\to g$ uniformly on $[c,d]$.
@@
denote the mesh norm of the irregular grids. Then the Taylor-accelerated neural network interpolation operator $\widetilde{T}_{n,r}$ satisfies
@@
Assume that the sequence of irregular grids satisfies the refinement condition: there exists a constant $C>0$, independent of $n$, such that
@@
Such a condition is fulfilled, for instance, by any quasi-uniform family of grids of $[c,d]$.
@@
Although the operator introduced in [7] is defined for all continuous functions $g\in C[c,d]$, its higher-order accuracy is obtained through a Lagrange-type reconstruction involving $(r+1)$ neighboring sampling nodes. On irregular grids, this may lead to additional numerical complexity and reduced stability due to the behaviour of Lagrange basis polynomials.
@@
2.2 Multivariate Extension
We now extend the above construction to the multivariate setting. Let $d\in\mathbb{N}$ and consider the $d$-dimensional box domain
$\Omega:=[a_{1},b_{1}]\times\cdots\times[a_{d},b_{d}]\subset\mathbb{R}^{d}.$
Let $\eta\in D(m)$ and let $\sigma:\mathbb{R}\to\mathbb{R}$ be the associated activation function introduced in Section 1. We define the multidimensional activation function
$\Psi_{\sigma}:\mathbb{R}^{d}\to\mathbb{R}$
by
$\Psi_{\sigma}(y):=\prod_{i=1}^{d}\sigma(y_{i}),\qquad y=(y_{1},\dots,y_{d})\in\mathbb{R}^{d}.$
The function $\Psi_{\sigma}$ is nonnegative, compactly supported, and satisfies
$\|\Psi_{\sigma}\|_{\infty}:=\sup_{y\in\Omega}|\Psi_{\sigma}(y)|\leq 1.$
Multivariate irregular grids. For each coordinate direction $i\in\{1,\dots,d\}$, we consider an irregular partition of the interval $a_{i},b_{i}]$ given by
$a_{i}=z_{i0}<z_{i1}<\cdots<z_{in}=b_{i}.$
We assume that there exist constants $0<d_{1}\leq d_{2}$, with $d_{2}<2d_{1}$, such that
$d_{1}\leq z_{i,k+1}-z_{ik}\leq d_{2},\qquad k=0,1,\dots,n-1,\ \ i=1,\dots,d.$
For a multi-index $k=(k_{1},\dots,k_{d})\in\{0,\dots,n\}^{d}$, we define the corresponding grid node
$z_{k}:=(z_{1k_{1}},\dots,z_{dk_{d}})\in\Omega.$
As in the one-dimensional case, to ensure consistency near the boundary and preserve the interpolation property, the nodes are extended by setting
$z_{i,n+\ell}:=z_{i,n-r+\ell-1},\qquad\ell=1,\dots,r,\ \ i=1,\dots,d.$
Multivariate Taylor polynomial. Let $g\in C^{r}(\Omega)$. For each node $z_{k}$, we denote by $P_{r}(g;y,z_{k})$ the multivariate Taylor polynomial of total degree $r$ of $g$ centered at $z_{k}$, defined by
$P_{r}(g;y,z_{k})=\sum_{|\alpha|\leq r}\frac{D^{\alpha}g(z_{k})}{\alpha!}(y-z_{k})^{\alpha},\qquad y\in\Omega,$
where $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index, $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}$, $\alpha!=\alpha_{1}!\cdots\alpha_{d}!$, and
$D^{\alpha}g:=\frac{\partial^{|\alpha|}g}{\partial y_{1}^{\alpha_{1}}\cdots\partial y_{d}^{\alpha_{d}}}.$
Definition 2.11.
Let $g\in C^{r}(\Omega)$. Then multivariate Taylor-accelerated neural network interpolation operator is defined by
$\widetilde{T}_{n,r}^{(d)}(g,y)=\frac{\displaystyle\sum_{k\in\{0,\dots,n\}^{d}}\Psi_{\sigma}\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)P_{r}(g;y,z_{k})}{\displaystyle\sum_{k\in\{0,\dots,n\}^{d}}\Psi_{\sigma}\!\left(\frac{2m}{d_{1}}(y-z_{k})\right)},\qquad y\in\Omega.$
The extension of the above construction and approximation results from the univariate setting to $\mathbb{R}^{d}$ does not introduce additional analytical difficulties. Since $\mathbb{R}^{d}$ is finite dimensional, all norms are equivalent and the properties of the activation function $\Psi_{\sigma}$ are preserved under the product structure. Consequently, the proofs of convergence, interpolation, and quantitative error estimates follow the same arguments as in the one-dimensional case, with only notational changes and dimension-dependent constants. This situation stands in sharp contrast with approximation problems posed in infinite-dimensional spaces, where genuinely new analytical phenomena typically arise.
@@ 3.1 Activation Function and Grid Construction
All experiments were implemented in MATLAB. The sampling nodes were chosen as nested quasi-uniform irregular grids of $[0,1]$, generated by applying small random perturbations to a uniform grid while preserving the mesh regularity condition $\frac{d_{2}}{d_{1}}<2.$ This construction ensures that the theoretical assumptions are satisfied and allows us to study convergence as the number of nodes $n$ increases.
@@ 3.2 Error Convergence and Comparison
where $T^{\mathrm{Lag}}_{n,r}$ is the Lagrange-based neural network interpolation operator on irregular grids introduced by Costarelli et al.[7].
@@
To highlight the improvement with respect to the Lagrange-based construction, Figures 2-4 compare the Taylor and Lagrange interpolation operators for each fixed order $r=1,2,3$. In all cases, the Taylor-accelerated interpolation operator provides smaller errors, demonstrating the effectiveness of the local Taylor enrichment.
Figure 2: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=1$.
Figure 3: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=2$.
Figure 4: Error comparison between Taylor and Lagrange neural network interpolation operators for $r=3$.
@@ 4 Conclusion
We proved that the operators are well defined, uniformly bounded, and preserve exact interpolation at the grid nodes. Moreover, they reproduce polynomials up to the prescribed order and satisfy direct approximation estimates in terms of higher-order moduli of smoothness, leading to convergence rates of order $r+1$ for sufficiently smooth functions. In particular, under asymptotically quasi-uniform grids, the error decays as
@@
Numerical experiments confirmed the theoretical findings and showed that the Taylor-accelerated operator provides improved accuracy compared not only with classical first-order neural network interpolation operators, but also with the recently proposed Lagrange-based neural network interpolation operator on irregular grids. This highlights the advantage of incorporating local Taylor information in achieving higher-order convergence on non-uniform sampling nodes.
Future research directions include adaptive irregular meshes and the development of stochastic or fuzzy variants of the proposed Taylor-accelerated framework.